OpenStudy (anonymous):

A video game sets the points needed to reach the next level based on the function g(x) = 10(2)x − 1, where x is the current level. The hardest setting promises to multiply the points needed in each level according to the function h(x) = 2x. How many points will a player need on the hardest setting of level 5? 10 points 160 points 1600 points 16000 points

OpenStudy (tkhunny):

Unfortunately, your g(x) doesn't make any sense. Figure out how to write it more clearly.

OpenStudy (anonymous):

g(x)=10(2)^x-1

OpenStudy (tkhunny):

\(g(x) = 10\cdot 2^{x-1}\)?

OpenStudy (anonymous):

yes

OpenStudy (tkhunny):

\(g(1) = 10\cdot 2^{1-1} = 10\cdot 2^{0} = 10\cdot 1 = 10\) \(g(2) = 10\cdot 2^{2-1} = 10\cdot 2^{1} = 10\cdot 2 = 20\) \(g(3) = 10\cdot 2^{3-1} = 10\cdot 2^{2} = 10\cdot 4 = 40\) I'm not convinced. Doesn't seem to check out.

OpenStudy (anonymous):

im confused @tkhunny